40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12552     Entry time: Wed Oct 12 13:34:28 2016     In reply to: 12551
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM LR problematic again 

Looks like what were PRM problems are now seen in the SRM channels, while PRM itself seems well behaved. This supports the hypothesis that the satellite box is problematic, rather than any in-vacuum shenanigans.

Eric noted in this elog that when this problem was first noticed, switching Satellite boxes didn't seem to fix the problem. I think that the original problem was that the Al foil shorted the contacts on the back of the OSEM. Presumably, running the current driver with (close to) 0 load over 2 months damaged that part of the Satellite box circuitry, which lead to the subsequent observations of glitchy behaviour after the pumpdown. Which begs the question - what is the quick fix? Do we try swapping out the LM6321 in the LR LED current driver stage? 

GV Edit Nov 2 2016: According to Rana, the load of the high speed current buffer LM6321 is 20 ohms (13 from the coil, and 7 from the wires between the Sat. Box and the coil). So, while the Al foil was shorting the coil, the buffer would still have seen at least 7 ohms of load resistance, not quite a short circuit. Moreover, the schematic suggests that that the kind of overvoltage protection scheme suggested in page 6 on the LM6321 datasheet has been employed. So it is becoming harder to believe that the problem lies with the output buffer. In any case, we have procured 20 of these discontinued ICs for debugging should we need them, and Steve is looking to buy some more. Ben Abbot will come by later in the afternoon to try and help us debug.

ELOG V3.1.3-